# MLM: Learning Multi-task Loco-Manipulation Whole-Body Control for Quadruped Robot with Arm

### v1

@@ IV-C Real-world Experiments
We further validate the feasibility and effectiveness of our method through real-world experiments. The experiments include both teleoperation via a handheld gripper and DP-based inference execution. All tasks were accomplished using a unified trained whole-body policy. Our approach demonstrates excellent performance in handling targets across all scenarios, achieving zero-shot transfer from simulation to the real world. Please refer to our videos in the supplemental materials.
Handheld Gripper Teleoperation: Figure 1 I–V showcase several real-world demonstrations of whole-body loco-manipulation using the teleoperation, such as cleaning up paper scraps, picking up trash, storing toys, unplugging chargers, and removing bolts to open doors. Figure 1 III demonstrates a long-range task. It can be observed that when performing ground-level tasks, such as picking up trash, the quadruped robot assists the arm by leaning downward, enabling the end-effector to reach lower positions. Conversely, tasks well above body height require lifting the body. When the target position lies well beyond the current workspace, the quadruped robot exploits its locomotion capability to reposition itself, thereby reaching the distant site. These behaviors demonstrate exceptional whole-body coordination capabilities. We also evaluated the tasks for training and recorded the tracking performance in Table IV. The results show that the real-world deployment achieves tracking performance comparable to that of the simulation.
TABLE IV: Real-world experiment results. (cm/radian/%)
@@
Inference through DP: As shown in Figure 1 VI–VII, we demonstrate the method’s ability to autonomously execute tasks through a pre-trained DP. Specifically, we input the information from a GoPro camera mounted on the robotic arm’s wrist, along with the arm’s states, into the DP. The DP then autonomously generates the trajectory of the target object. The RL whole-body policy is then used to control the joints for execution. Its performance is mainly limited by the results of DP inference. The experimental results indicate that our method can leverage the DP to plan and autonomously perform multi-task operations, as shown in Table IV.
@@ V Conclusion
We proposed MLM, a reinforcement learning framework that leveraged both real-world and simulation data to address multi-task whole-body loco-manipulation control for quadruped robots. To balance performance while learning multiple task trajectories, we integrate a data library with an adaptive sampling mechanism. To address the scenario where only historical trajectories can be observed during teleoperation, we propose a Trajectory-Velocity Prediction policy network to predict future trajectories, thereby improving policy execution performance. Velocity estimation supports task execution across different spatial ranges. Finally, simulations and real-world experiments on Go2 and Airbot Play robot highlight the good performance of the proposed method. In fact, our approach can be extended to enable quadruped robots to perform more tasks in new environments, fully utilizing their locomotion and manipulation capabilities. This depends on higher-level decision-making and planning. Future integration of vision-language models and navigation could address this limitation.

### v2

@@ IV-C Real-world Experiments
We further validate the feasibility and effectiveness of our method through real-world experiments, including teleoperation and DP inference. All tasks are executed with a unified whole-body policy, showing strong performance across tasks and enabling zero-shot transfer from simulation to reality. Please refer to our videos in the supplemental materials.
Handheld gripper teleoperation: Figure 1 I–V showcase real-world demonstrations of whole-body loco-manipulation via teleoperation, including cleaning papers, picking up trash, storing a toy, unplugging a charger, and removing a bolt to open a door. Figure 1 III demonstrates a long-range task. For ground-level tasks, such as picking up trash, the quadruped robot assists the arm by leaning downward, enabling the end-effector to reach lower positions. Tasks well above body height require lifting the body. When the target position lies beyond the current workspace, the quadruped robot leverages its locomotion to reposition itself. These behaviors demonstrate exceptional whole-body coordination capabilities. Two tasks in Table V show that the real-world deployment achieves tracking performance comparable to that of the simulation.
TABLE V: Real-world experiment results. (cm / radian / %)
@@
Inference through DP: As shown in Figure 1 VI–VII, we demonstrate the ability to autonomously execute tasks through a pre-trained DP. Specifically, we input the information from a GoPro camera mounted on the arm’s wrist, along with the arm’s states, into the DP. The DP then generates the trajectory to the target object. The RL whole-body policy is then used to control the joints for execution. Its performance is mainly limited by the results of DP inference. The experimental results indicate that our method can leverage the DP to plan and autonomously perform multi-task operations, as shown in Table V.
@@ V Conclusion
We proposed MLM, a reinforcement learning framework that leveraged both real-world and simulation data to address multi-task whole-body loco-manipulation control for quadruped robots. To balance performance across multiple tasks, we integrate a data library with an adaptive sampling mechanism. For teleoperation where only historical trajectories are observable, we introduce a Trajectory-Velocity Prediction policy network to predict future trajectories, enhancing execution performance. Velocity estimation supports task execution across different spatial ranges. Simulations and real-world experiments on Go2 and Airbot Play also demonstrate the effectiveness of the proposed method. Looking ahead, our method can be extended to more tasks in new environments via higher-level decision-making (e.g., vision–language models). Moreover, handling highly dynamic tasks may benefit from faster sensor update rates and better actuation response capabilities.